how does the 4th root of 64 simplify into 2 times the 4th root of 4? please show me how to work it out...
Answer by lwsshak3(11628)

how does the 4th root of 64 simplify into 2 times the 4th root of 4? please show me how to work it out
**
64^1/4=(2^6)^1/4=2^6/4=2^3/2=√8=2√2
2*4^1/4=2*(2^2)^1/4=2*2^2/4=2*2^1/2=2√2